# Gotta Hear Them All: Sound Source Aware Vision to Audio Generation

### v3

@@ A.4.3 Ablation of Remixer Architecture
The optimal setup of the Sound Source Remixer’s architecture is found through ablations recorded in Tab. 11. Increasing the number of attention layers slightly increases single-source generation performance. However, the multi-source generation quality is significantly sacrificed as attention layers stack deeper. Since our work’s primary focus is to tackle the multi-sound-source V2A generation problem, we choose the one-attention-layer architecture as the optimal setting for the Sound Source Remixer. We use the same architecture for the Temporal Aggregation module.
Appendix A.5 Setup of Subjective Evaluation
@@
We disseminate an online survey for the subjective evaluation and collect results from 20 participants to measure generation fidelity and relevance of our method along with baselines. The baseline methods include Im2Wav, Diff-Foley-Image, S&H-Text, and V2A-Mapper as described in Appendix A.1. In the first survey section, we ask the participants to sign a consent form as illustrated in Fig. 12 (a). The non-consenting participants are screened out without any data collection. In the second section, we ask 20 fidelity-rating questions without visual cues following Fig. 12 (b). To unify the comparison context, we include a short tag in the question describing the ground-truth audio content. For each generated sample, the testee is asked to give out fidelity rating on a 1-5 scale. In the third section, we ask 20 relevance-rating questions given the visual condition used during generation, which is depicted in Fig. 12 (c). The ratings are also collected on a 1-5 scale.
@@
After data collection, we thoroughly anonymize the participant information to deidentify any personal data. We then compute the Mean Opinion Score (MOS) [51] respectively from the fidelity and relevance ratings.
Appendix A.6 Why Not Cascaded Composition?
@@
The most straightforward way to composite multimodal sound sources into a single audio is to generate an audio track for each source condition via video-to-audio or text-to-audio models and overlay them together. However, such a cascading audio synthesis system lacks interaction, context and style awareness when integrating multiple sound sources, which are keys to a convincing audio scene. We show with qualitative examples in this section that our SSV2A composition achieves these features. These examples are best experienced on our website at the Composition Comparisons section.
A.6.1 Interaction Awareness
@@
We generate a drum-only audio clip and a bass-only clip with SSV2A. Simply overlaying these clips yields a mixed track with no interactions between these instruments, which is shown in Fig. 9. With our SSV2A composition, we are able to generate an appealing piece of drum-bass music with rich interactions.
A.6.2 Context Awareness
@@
We generate audio clips of a normal speech and an academic conference in Fig. 10. Cascaded composition synthesizes an audio clip with conflicting talkers and no conference room reverb. Our SSV2A composition properly transfers the speech style into a reverberated academic presentation.
A.6.3 Style Awareness
@@
In Fig. 11, we generate an audio clip of a talking man and another clip of police activities. The cascaded composition yields chaotic sounds as the police event context is not perceived. Our SSV2A composition successfully picks up this global cue and generates a police officer’s voice followed by a gun loading/shooting sound to indicate police events.
@@
Figure 13: Screenshot of web demo application. Please visit our website and click the demo button to access this interface.
Appendix A.7 Future Improvements on Limitations
A.7.1 Temporal Synchronization
@@
Our simple TA module has shown decent synchronization capability by attending to global audio scene semantics, as evidenced by WMAO evaluations in the Tab. 3. In reality, each sound source can have different temporal “activation” intervals in an audio scene. For example, a dog may bark in only the first three seconds of a video, followed by a baby laughing. In future works, we propose to perceive each sound source’s activations locally in time, similar to the tracklet detection [10] concept in video multi-object tracking. With this finer-grained temporal aggregation approach, we aim to further enhance SSV2A’s temporal alignment performance.
A.7.2 Sensitivity of Audio Conditioning
@@
As stated in the limitations section, the lower sensitivity of our method’s audio conditioning compared to other modalities’ input is due to the lack of CLIP semantics for audio sound source inputs. An existing method, Wav2CLIP [63], translates CLAP audio embeddings to highly relevant CLIP embeddings. However, it operates on ViT-B/32 instead of ViT-L/14, which is the CLIP variant SSV2A employs. In future works, we plan to train a Wav2CLIP model compatible with SSV2A to address the audio sensitivity issue.
Appendix A.8 Implementation
We upload the implementation and pretrained weights of SSV2A to github at https://github.com/wguo86/SSV2A. Please follow the readme instructions to set up a virtual environment and execute the designated scripts.
Appendix A.9 Web Demo Instructions
We implement a web application for interactive demonstration, as illustrated in Fig. 13. The left panel supports entries of generation conditions from image, text and audio. The right panel offers further hyperparameter tuning. We host this application on our own machine via a secure internet tunnel, strictly avoiding any collection of user data.
Appendix A.10 Ethical Statement
@@
Our human evaluation is strictly anonymized without collecting any sensitive personal data. We also obtain explicit verbal consent from participants by asking them to sign a data collection agreement form before survey and screening out non-consenting participants. We intend to make our curated dataset, VGGS3, publicly available to contribute to the visual-audio research community. Our V2A method complements videos and images with convincing audio tracks. Its application may have malicious outcomes in deepfake [62] multimedia products if used without censorship. Multiple multimedia deepfake detection approaches have been proposed including audio deepfake detection [52]. We are committed to contribute ample generation samples for strengthening the learning of these detectors.

### v4

@@ 4.3 Ablation of Remixer Architecture
The optimal setup of the Sound Source Remixer’s architecture is found through ablations recorded in Tab. 10. Increasing the number of attention layers slightly increases single-source generation performance. However, the multi-source generation quality is significantly sacrificed as attention layers stack deeper. Since our work’s primary focus is to tackle the multi-sound-source audio generation problem, we choose the one-attention-layer architecture as the optimal setting for the Sound Source Remixer. We use the same architecture for the Temporal Aggregation module.
5 More Experiments
Method VGG-$S$ MUSIC
[…]
Single-Source
Diff-Foley 7.212 39.309 11.045 4.099 27.633 79.068 9.286 5.899
Frieren 2.799 28.340 11.134 4.326 9.001 56.014 12.613 6.650
MultiFoley 3.890 30.244 11.827 5.648 - - - -
MMAudio 2.882 20.659 13.971 5.991 3.471 33.339 14.662 7.866
SS2A (Ours) 2.270 14.433 12.176 5.336 8.353 25.393 13.392 7.182
Multi-Source
Diff-Foley 13.373 75.829 12.209 4.789 12.423 105.299 8.561 4.843
Frieren 10.601 66.896 11.884 4.526 4.803 69.921 11.857 5.979
MultiFoley 9.082 67.201 12.648 5.621 - - - -
MMAudio 8.178 50.657 13.425 6.053 5.236 31.880 13.149 6.693
SS2A (Ours) 6.837 45.647 12.195 5.658 3.562 30.260 12.436 6.021
Table 11: Video to audio comparisons. The first and second places are bolded and underlined, respectively.
Method WMAO $\downarrow$
Top-1 Top-5
Diff-Foley 1.227 1.127
FoleyCrafter 1.247 1.167
V2A-Mapper 1.299 1.197
Ours w/o TA 1.292 1.205
Ours 1.243 1.172
Table 12: AVSync synchronization tests. Bold, underline, and italic mark first, second, and third placements.
5.1 Video-to-Audio Objective Tests
We employ the same metrics in image-to-audio tasks to judge the generation quality of video-to-audio samples. Our method performs competitively in Tab. 11, especially in multi-source scenarios. MultiFoley isn’t open-sourced, so MUSIC tests are void. We lead in fidelity while MMAudio leads in relevance, likely due to its joint audio-visual (AV) and audio-text (AT) training. Our Remixer, trained only on AV, aligns more with AV distributions (fidelity), while MMAudio benefits from semantics in AT (relevance).
5.2 Video-to-Audio Synchronization Tests
Weighted Mean Absolute Offset.
To prove the efficacy of our TA mechanism for video-audio synchronization, we employ SynchFormer (Iashin et al. 2024) to predict the temporal Weighted Mean Abolute Offset (WMAO) in seconds between the original video and the generated audio on the AVSync15 (Zhang et al. 2024a) dataset. We weight and sum the top-k predictions of SyncFormer with their confidence scores. A lower WMAO perceives smaller drifts between video-audio signals, indicating higher synchronization.
Our method generates competitive video-audio synchronization as shown in Tab. 12. The ablation in TA mechanism shows that it is key to SS2A’s temporal alignment capability. Moreover, the results indicate that our nonlinear TA function performs better than V2A-Mapper’s linear setup.
6 Setup of Subjective Evaluation
@@
We disseminate an online survey for the subjective evaluation and collect results from 20 participants to measure generation fidelity and relevance of our method along with baselines. The baseline methods include Im2Wav, S&H-Text, and V2A-Mapper as described in Sec. 1. In the first survey section, we ask the participants to sign a consent form as illustrated in Fig. 12 (a). The non-consenting participants are screened out without any data collection. In the second section, we ask 20 fidelity-rating questions without visual cues following Fig. 12 (b). To unify the comparison context, we include a short tag in the question describing the ground-truth audio content. For each generated sample, the testee is asked to give out fidelity rating on a 1-5 scale. In the third section, we ask 20 relevance-rating questions given the visual condition used during generation, which is depicted in Fig. 12 (c). The ratings are also collected on a 1-5 scale.
@@
After data collection, we thoroughly anonymize the participant information to deidentify any personal data. We then compute the Mean Opinion Score (MOS) (Sector 1996) respectively from the fidelity and relevance ratings.
7 Why Not Cascaded Composition?
@@
The most straightforward way to composite multimodal sound sources into a single audio is to generate an audio track for each source condition via video-to-audio or text-to-audio models and overlay them together. However, such a cascading audio synthesis system lacks interaction, context and style awareness when integrating multiple sound sources, which are keys to a convincing audio scene. We show with qualitative examples in this section that our SS2A composition achieves these features. These examples are best experienced on our website at the Composition Comparisons section.
@@
7.1 Interaction Awareness
We generate a drum-only audio clip and a bass-only clip with SS2A. Simply overlaying them yields a mixed track with no interactions between these instruments, as shown in Fig. 9. With SS2A composition, we are able to generate an appealing piece of drum-bass music with rich interactions.
7.2 Context Awareness
We generate audio clips of a normal speech and an academic conference in Fig. 10. Cascaded composition synthesizes an audio clip with conflicting talkers and no conference room reverb. Our SS2A composition properly transfers the speech style into a reverberated academic presentation.
7.3 Style Awareness
In Fig. 11, we generate an audio clip of a talking man and another clip of police activities. The cascaded composition yields chaotic sounds as the police event context is not perceived. Our SS2A composition successfully picks up this global cue and generates a police officer’s voice followed by a gun loading/shooting sound to indicate police events.
8 Limitations and Future Improvements
Two limitations exist in SS2A. First, we address the video-to-audio synchronization in SS2A with a naive temporal module. Existing works (Iashin et al. 2022, 2024) show that temporal alignment is a nontrivial problem due to the sparsity of synchronization cues in both time and space. Second, we observe that SS2A is less sensitive to audio conditioning than visual or text inputs. We suspect that this phenomenon is due to the lack of CLIP semantics when the Sound Source Remixer is prompted with audio conditions. We propose solutions to these issues as future works in this section.
8.1 Temporal Synchronization
@@
Our simple TA module has shown decent synchronization capability by attending to global audio scene semantics, as evidenced by WMAO evaluations in the Tab. 3. In reality, each sound source can have different temporal “activation” intervals in an audio scene. For example, a dog may bark in only the first three seconds of a video, followed by a baby laughing. In future works, we propose to perceive each sound source’s activations locally in time, similar to the tracklet detection (Ciaparrone et al. 2020) concept in video multi-object tracking. With this finer-grained temporal aggregation approach, we aim to further enhance SS2A’s temporal alignment performance.
8.2 Sensitivity of Audio Conditioning
@@
As stated in the limitations section, the lower sensitivity of our method’s audio conditioning compared to other modalities’ input is due to the lack of CLIP semantics for audio sound source inputs. An existing method, Wav2CLIP (Wu et al. 2022), translates CLAP audio embeddings to relevant CLIP embeddings. However, it operates on ViT-B/32 instead of ViT-L/14, which is the CLIP variant SS2A employs. In future works, we plan to train a Wav2CLIP model compatible with SS2A to address the audio sensitivity issue.
9 Ethical Statement
@@
Our human evaluation is strictly anonymized without collecting any sensitive personal data. We also obtain explicit verbal consent from participants by asking them to sign a data collection agreement form before survey and screening out non-consenting participants. We intend to make our curated dataset, VGGS3, publicly available to contribute to the visual-audio research community. Our method complements videos and images with convincing audio tracks. Its application may have malicious outcomes in deepfake multimedia products if used without censorship. Multiple multimedia deepfake detection approaches have been proposed including audio deepfake detection (Shaaban, Yildirim, and Alguttar 2023). We are committed to contribute generation samples for strengthening the learning of these detectors.